OpenStudy (anonymous):

Factor Completely: 12a3d2 − 6ad3. Awarding medals!

OpenStudy (ivettef365):

now you can get the GCF, as follows: between 12 and 6 = 6 " a^3 and a = a (the smallest always) " d^2 and d^3 = d^2 ( the smallest)

OpenStudy (anonymous):

So the GCF is 6ad^2

OpenStudy (anonymous):

Correct. So write the GCF, and then next to it, a single set of parenthesis. The parenthesis should have as many terms as the thing you're pulling the GCF out of, so in this case, it will have 2. \(\Large 6ad^2(\text{_ } \pm \text{ _})\)

OpenStudy (anonymous):

Now your job is to fill in those spots in the parenthesis so that when you distribute the GCF in, the result is what you started with.

OpenStudy (anonymous):

Ask yourself, "What can I write here so that if I multiplied the GCF in, I would get ______?"